OpenStudy (anonymous):

Just a quick answer check 5(14-2)ˆ2 Divided by 2 A.60 B.30 C.72 D.360 I got 60 for my answer :D

OpenStudy (anonymous):

sorry lol. 14-2=?

OpenStudy (anonymous):

here let me copy and paste the exact thing

OpenStudy (anonymous):

14-2=12 5(12)*2

OpenStudy (anonymous):

5(14 - 2)2 ÷ 2

OpenStudy (anonymous):

What is 5 times 12 @AnimeUnity

OpenStudy (anonymous):

its 60

OpenStudy (anonymous):

im unsure of if I'm right or not i got 360 as one answer and 60 as my next ;-; but I'm going with 60 cuz it seems more realistic

OpenStudy (anonymous):

http://web2.0calc.com/

OpenStudy (anonymous):

60*2=120/2=60

OpenStudy (anonymous):

remember your order of operations :)

OpenStudy (anonymous):

So am i right :)

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@Tazmaniadevil Please explain or show your work and not just state the answer. Thx

OpenStudy (anonymous):

Cool thanks guys ;) this websites a big help with everything you guys rock

OpenStudy (anonymous):

i did i showed how i got it i used http://web2.0calc.com/

OpenStudy (anonymous):

After you gave the answer

OpenStudy (anonymous):

Taz thanks man :D

OpenStudy (anonymous):

shh i have been hurry before you this is my third year here check my stats ik what im doing

OpenStudy (anonymous):

np i feel the struggle of math im on pre calc 12 grade haha

OpenStudy (anonymous):

xD yea dude i try to get my asseinments turned in before the actual day hits I'm already on thursdays work

OpenStudy (anonymous):

i im on next weeks work lol

OpenStudy (anonymous):

Noice i guess we both read survival guides to school lol

OpenStudy (anonymous):

At least i do xD

OpenStudy (anonymous):

ik lol hahah only thing i can do english 4 is my last class im so happy lol

OpenStudy (anonymous):

yea well hey man thanks for the help i got to continue to the next day :D thanks a lot

OpenStudy (anonymous):

Everyone

OpenStudy (anonymous):

aight seya later

OpenStudy (anonymous):

have a good day

OpenStudy (anonymous):

you to man